OpenStudy (anonymous):

Medal You bought a car for $7000 down and made payments of $699.50 each month for 36 months. (a) Find the amount of the payments over the 36 months. $ (b) Find the total cost of the car. $

OpenStudy (anonymous):

a=25,182 b=32182

OpenStudy (anonymous):

how did you get this

OpenStudy (anonymous):

i simply multiplied 699.50 times the 36 months for the first answer an for the second i just added the first number to the 7,000 down payment

OpenStudy (tkhunny):

Maybe 7000 + 36*699.50 = 7000 + 25,182 = 32,182 -- Depends what it means by "a payment". The second question may be impossible. Does it mean the "Price". We don't know that.

OpenStudy (anonymous):

ok thank you